OpenStudy (anonymous):

4x^3-25x and 4x^3-7x^2-x they're both factorization equations, how do you solve them?

OpenStudy (anonymous):

4x^3-25x=x(4x^2-25)=0

OpenStudy (anonymous):

4x^3-7x^2-x

OpenStudy (anonymous):

do the same

OpenStudy (anonymous):

\[4x^3 - 25x = {x}(4x^2 - 25) = x({(2x)^2-(5^2))}=x(2x+5)(2x-5)\]

OpenStudy (anonymous):

so now we need to factorize Actually whenever if any equation ends at x that means there is a root x=0 so you have to take x common and then put the other part =0

OpenStudy (anonymous):

\[4x^3 - 7x^2 -x = {x}(4x^2 - 7x-1)\]